OpenStudy (anonymous):

Translate the words into an appropriate function. Express the area, A, of a circle as a function of its diameter d.

Directrix (directrix):

Area circle = pi * r^2 The diameter of a circle is twice the radius of the circle.

Directrix (directrix):

So, r = d/2 where r is the radius and d is the diameter

Directrix (directrix):

Area circle = pi * r^2 Area circle = pi * (d/2)^2 @maggielucci What is d/2 times d/2 = ?

Directrix (directrix):

The idea is to simplify this: (d/2)^2